OpenStudy (anonymous):

A student is running a 10-kilometer race. He runs 1 kilometer every 4 minutes. Select the function that describes his distance from the finish line.

OpenStudy (anonymous):

f(x) = 1 over 4x + 4 f(x) = - 1 over 10x + 4 f(x) = 1 over 10x + 10 f(x) = - 1 over 4x + 10